OpenStudy (anonymous):

2x+y=4 3x-2y=-15, solve by substitution. can someone show me step by step how to work this please

OpenStudy (anonymous):

Sure :) start by solving the first equation for y: 2x+y=4 -2x -2x y=4-2x Then substitute this into the second equation: 3x-2y=-15 3x-2(4-2x)=-15 Now solve for x: 3x-8--4x=-15 3x-8+4x=-15 7x-8=-15 7x=-7 x=-1 Now substitute this into y=4-2x to find y: y=4-2(-1) y=4--2 y=4+2 y=6 so x=-1 and y=6 Hope this helped!

OpenStudy (anonymous):

awh! thank you so much! :)

OpenStudy (anonymous):

You're welcome!